A pension fund manager is considering three mutual funds. The first is a stock fund, the second is a long-term government and corporate bond fund, and the third is a T-bill money market fund that yields a rate of 7%. The probability distribution of the risky funds is as follows:

Expected Return Standard Deviation
Stock fund (S) 16 % 38 %
Bond fund (B) 12 21

The correlation between the fund returns is 0.12.

You require that your portfolio yield an expected return of 11%, and that it be efficient, on the best feasible CAL.

a. What is the standard deviation of your portfolio? (Round your answer to 2 decimal places.)

b. What is the proportion invested in the T-bill fund and each of the two risky funds? (Round your answers to 2 decimal places.)

Proportion Invested in:

T-bill funds = ____ %

Stocks = ____ %

Bonds = ____ %
Answer #1

First we need to find the weight of stock and bond within in the risky portfolio through the equation of the optimal portfolio:

(r Stock – Frisk free)o Bond - (r Bond - Trisk free)p(Stock X Bond) W Stock = 7 (rStock - Trisk free) Bond + (r Bond - Prisk

W Bond = 1-W Stock WBond = 1-0.35 W Bond = 0.65 or 65%

Now we need to calculate the expected return and the standard deviation of this risky portfolio:

Risky Portfolio Expected return is calcualted by solving the following equation:


Expected retur n risky portfolio = Weightstockr Stock + Weight Bond Bond Expected retur risky portfolio = 0.35 x 16 + 0.65 x

Risky Portfolio standard deviation is calculated by solving the following equation:


Orisky portfolio = V(Weight stock stock)+(Weight Bond Bond) +2Weight stock stock Weight Bond Bond X correlartion stock, Bond Orisky portfolio = V(0.35x38)2+(0.65 x 212+2(0.35 x 38 x0.65 x 21 x0.12)

Orisky portfolio = 20.17%

Now we need to combine this risky portfolio with the tbill to generate the expected return of 11%

Portfolio Expected return is calcualted by solving the following equation:
Expected return portfolio = Weight Riskyr Risky + WeightTbilrtbill + 11% = Weight Risky X 13.4+ (1 - Weight Risky) x 7 Weight

Now as the Tbill has no risk so the entire risk comes from the risky portfolio so that is calculated is as follows:

Oportfolio = Weight Risky timesorisky Oportfolio = 0.625 times20.17 portfolio = 12.61%

Below is the calcualtion of each asset:

Asset Portfolio weight Asset weight Final weights = Portfolio Weigh x Asset Weight
Tbill 37.50% 100.00% 37.50%
Stock 62.50% 35.00% 21.88%
Bond 62.50% 65.00% 40.63%